OpenStudy (anonymous):

The position function for an object is given by s(t)=6t^2+240t, where s is measured in feet and t is measured in seconds. Find the velocity of the object when t=2 seconds.

OpenStudy (anonymous):

Velocity is the derivative of the position with respect to time. Are you given the position with respect to time. All you have to do is differentiate.

OpenStudy (anonymous):

Do I have to plug in the 2?

OpenStudy (anonymous):

v(t)=ds/dt v(t)=12t+240 v(2)=12*2+240=264 feet/second

OpenStudy (anonymous):

You plug the 2 into the derivative.

OpenStudy (anonymous):

You mean you have to use: \[ \sin(x/2) = \pm\sqrt{\frac{1-\cos(x)}{2}} \]?

OpenStudy (anonymous):

In that case split it into: 1 + 12

OpenStudy (anonymous):

you're going to have to split it regardless.

OpenStudy (anonymous):

just differentiate the given function with respect to 't' It will give you velocity as a function of 't'. just plug in the required value i.e '2' to obtain velocity at time 2 seconds...